OpenStudy (anonymous):

please final approaching!!!! tomorrow! find lim x-->6 f(x) if f(x)= x-6, x> or = to 6 (x-6)^3, x<6

OpenStudy (unklerhaukus):

\[\lim\limits _{x\rightarrow6} (x-6)=(6-6)=\cdots\]

OpenStudy (anonymous):

help me Tweety?

OpenStudy (anonymous):

somebody?!?

OpenStudy (unklerhaukus):

i don't understand what the second and third parts of the question are asking for

OpenStudy (anonymous):

me neither :(

OpenStudy (unklerhaukus):

you have an answer for the first part right?

OpenStudy (anonymous):